The lease agreement and related facts indicate the following:

a. Leased equipment had a retail cash selling price of $400,000. Its useful life was six years with no residual value.

b. The lease term was six years and the lessor paid $315,000 to acquire the equipment (thus, selling profit).

c. Lessor’s implicit rate when calculating annual lease payments was 8%.

d. Annual lease payments beginning January 1, 2021, the beginning of the lease, were $80,117.

e. Incremental costs of commissions for brokering the lease and consummating the completed lease transaction incurred by the lessor were $8,500.

Required:

1. & 2. Prepare the appropriate entries for the lessor to record the lease and the initial payment at its commencement and any entry(s) necessary at December 31, 2021, the fiscal year-end. (Round your intermediate and final answers to the nearest whole dollar amount. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

Implicit Rate is 8%

Yearly Instalment 80,117 for lease period of 6 Year.

Calculate PV of the payment as of 1/January/2021

1 Year = 80,117

2 Year = 80,117 x 1/1.08 = 74,182

3 Year = 80,117 x 1/(1.08)2 = 68,687

4 Year = 80,117 x 1/(1.08)3 = 63,599

5 Year = 80,117 x 1/(1.08)4 = 58,888

6 Year = 80,117 x 1/(1.08)5 = 54,526

So Total PV is 400,000

Journal Entry in the books of Lessor.

On 1/January/2021 - Lease Receivable Debit 400,000

To Assets Credit 400,000

On 1/January/2021 - Cash Debut 80,117

To Lease Receivable Credit 73,708

(80117 - 6409)

To Finance Income Credit 6,409

(80117 x 8/100)

On 31/December/2021- Lease Receivable 8,500

To Cash Credit 8,500